OpenStudy (babyslapmafro):

Could someone please verify if the following two functions are equivalent (click to see).

OpenStudy (babyslapmafro):

\[\frac{ \tan^4(4x)+2\tan^2(4x) }{ 16 }=\frac{ 1 }{ 16 }\sec^4(4x)\]

OpenStudy (babyslapmafro):

I was integrating and my answer is on the left, the answer in the book is on the right.

OpenStudy (babyslapmafro):

This is the integral \[\int\limits_{}^{}\tan(4x)\sec^4(4x)dx\]

OpenStudy (zehanz):

Doesn't look the same. First leave out the unimportant bits: set 4x =x and do not divide by 16. Then it would have to be that: \(\tan^4x+2\tan^2x=\sec^4x\)

OpenStudy (babyslapmafro):

\[u=\tan(4x)\] \[\int\limits_{}^{}(u)\sec^2(4x)(u^2+1)\frac{ du }{ 4\sec^2(4x) }\] \[=\frac{ 1 }{ 4 }\int\limits_{}^{}(u^3+u)du\]

OpenStudy (babyslapmafro):

\[=\frac{ 1 }{ 4 }[\frac{ u^4 }{ 4 }+\frac{ u^2 }{ 2 }]+C\]

OpenStudy (babyslapmafro):

\[=\frac{ \tan^4(4x) }{ 16 }+\frac{ \tan^2(4x) }{ 8 }+C\]

OpenStudy (babyslapmafro):

=(my original answer as-in my first post)

OpenStudy (babyslapmafro):

yay or nay

OpenStudy (zehanz):

u=tan(4x), so du=4sec(4x). You put in a factor du/(4sec(4x)), so that is 1. That is OK. You also have a factor u²+1 = tan²(4x)+1=sec²(4x). I go with that one too. Everything seems to be OK there!

OpenStudy (babyslapmafro):

ok so my answer is good? i tried punching both functions in my calc and got different answers, i was careful with the input.

OpenStudy (zehanz):

I would rather say: I do not see where you made a mistake... There is, somewhere. Below I made a graph of \(\tan^4(x)+2\tan^2(x)\) (blue) together with \(\sec^4(x)-1\) (red, dotted). So it seems your solution differs only a constant (-1) from the one in the book. Well, that is the end of the problem: you just need another integration constant!

OpenStudy (zehanz):

In other words, altough both solutions (without the integration constant!) are different, they only differ 1, so you have made no mistake after all. Congratulations!

OpenStudy (babyslapmafro):

ok thanks

OpenStudy (zehanz):

YW!